OpenStudy (anonymous):

What is the First Step In finding a Sum Or difference of two fractions with different denominators?

OpenStudy (anonymous):

Find the Least Common Denominator

OpenStudy (anonymous):

And Then?

OpenStudy (anonymous):

Here's an example

OpenStudy (anonymous):

\[\frac{ 2 }{ 8 }\] + \[\frac{ 4 }{ 12 }\] The least common denominator is 24, and 24/12=2, so multiply the 4 by 2. In 2/8, do 24/8=3. Then multiply 2 by 3.

OpenStudy (anonymous):

Then finally add or subtract.

OpenStudy (anonymous):

you add first Then add OR subtract?

OpenStudy (anonymous):

Oh nvm i got it

OpenStudy (anonymous):

Ok